How many ways can 7 students be assigned a seat in a classroom if there are 10 seats in a row?
Answer by Maths68(1474)

First Student can be assigned a seat = 10 ways
Since the 1 seat is occupied next student has 9 seats to sit
Second Student can be assigned a seat = 9 ways
Since the 2 seats are occupied next student has 8 seats to sit
Third Student can be assigned a seat = 8 ways
Since the 3 seats are occupied next student has 7 seats to sit
Fourth Student can be assigned a seat = 7 ways
Since the 4 seats are occupied next student has 6 seats to sit
Fifth Student can be assigned a seat = 6 ways
Since the 5 seats are occupied next student has 5 seats to sit
Sixth Student can be assigned a seat = 5 ways
Since the 6 seats are occupied next student has 4 seats to sit
Seventh Student can be assigned a seat = 4 ways
7 students can be assigned a seat in a classroom =10*9*8*7*5*4=604800 ways
It is a case of choosing 7 out of 10
nPr=n!/(n-r)!
10P7=10!/(10-7)!
10P7=10!/3!
10P7=10*8*7*6*5*4*3!/3!
Cancel out 3!
10P7=10*9*8*7*6*5*4=604800 ways